Compute the probability that a 5-card poker hand
a. is dealt to you that contains all hearts.
b. hand is dealt to you that contains four Aces.

Answer by math_tutor2020(3835)


Answers:
  1. 33/66640
  2. 1/54145
===============================================================================================

Work Shown:

Part (a)

13 hearts out of 52 cards total
13/52 = probability of getting a heart
12/51 = probability of getting a second heart
We do not put the cards back, aka no replacement.
Subtract 1 from numerator and denominator each time we take out another card.
11/50 = probability of getting a third heart
10/49 = probability of getting a fourth heart
9/48 = probability of getting a fifth heart

Multiply out the fractions
(13/52)*(12/51)*(11/50)*(10/49)*(9/48)
= (13*12*11*10*9)/(52*51*50*49*48)
= 154440/311875200
= 33/66640 is the probability of getting 5 hearts in a row.


Another approach:
n = 13 hearts
r = 5 slots to fill
Use the nCr combination formula. The order doesn't matter in a poker hand.
n C r = (n!)/(r!(n-r)!)
13 C 5 = (13!)/(5!*(13-5)!)
13 C 5 = (13!)/(5!*8!)
13 C 5 = (13*12*11*10*9*8!)/(5!*8!)
13 C 5 = (13*12*11*10*9)/(5!)
13 C 5 = (13*12*11*10*9)/(5*4*3*2*1)
13 C 5 = 154440/120
13 C 5 = 1287
There are 1287 ways to get a hand of nothing but hearts.
Follow similar steps to find 52C5 = 2598960 different five-card hands.
1287/2598960 = 33/66640

-----------------------------

Part (b)

There is 1 way to select the 4 aces where order doesn't matter.
There are 52-4 = 48 other cards that aren't an ace.
Ultimately there are 1*48 = 48 five-card hands that consist of 4 aces and some other card.

This is out of the 2598960 different five-card hands.
48/2598960 = 1/54145 represents the probability of getting four aces.

-----------------------------

For comparison purposes:
33/66640 = 0.000495 = 0.0495% 
 1/54145 = 0.000018 = 0.0018%
Each decimal value and percentage is approximate.

The event "getting 5 hearts" is more likely compared to the event "getting 4 aces".
Both events are really rare on their own.